OpenStudy (anonymous):

Solve the equation (4m-2)^2 - 2(4m-2)=15.

OpenStudy (lgbasallote):

expand first \[\implies (16m^2 - 16m + 4) - (8m - 4) = 15\] simplify \[\implies 16m^2 - 16m + 4 - 8m + 4 = 15\] combine like terms \[\implies 16m^2 -24m + 8 = 15\] subtract 15 from both sides \[\implies 16m^2 - 24m - 7 = 0\] use quadratic formula to solve for m

OpenStudy (anonymous):

Thanks you sir, I was stuck on that problem for 10 mins and i forgot to overlook factoring (4m-2)(4m-2).

OpenStudy (lgbasallote):

no worries. glad to help you out of the stump

OpenStudy (callisto):

A little bit late. But I did it in this way. (4m-2)^2 - 2(4m-2)=15 (4m-2)^2 - 2(4m-2) - 15=15 - 15 (4m-2)^2 - 2(4m-2)-15 = 0 Just consider (4m-2) as x, you'll get. x^2 - 2x -15 = 0 (x-5)(x+3) = 0 x= 5 or x=-3 Put back x=4m-2 into the solutions you got. 4m-2 = 5 or 4m-2 = -3 <- solve these two equations. If you understand how to do it, you can skip the step of letting x=4m-2...